# Evaluating Numerical Reasoning in Text-to-Image Models

### v1

@@ 1 Introduction
Using GeckoNum, we evaluate seven models chosen from three different model families (DALL·E 3, Imagen and Muse): we generate images for these models and collect human annotations to measure whether the images correctly match the prompts with respect to numerical reasoning. Our benchmark consists of text prompts, generated images, and a total of human annotations that we release: https://github.com/google-deepmind/geckonum_benchmark_t2i.
@@ 2 Related Work
Text-to-image benchmarking. Benchmarks used for evaluation of text-to-image models typically consist of a set of text prompts that target a specific capability. Some benchmarks, such as DALL·E 3 Eval [2] are more general and aim to capture real world use cases, while others contain a variety of challenges, such as prompts in DrawBench [21], PartiPrompts [32], and HEIM [15]. Others are developed to interrogate models on more specific capabilities such as alignment (e.g. TIFA [13], DSG [7] and Gecko [29]) or compositionality (e.g. T2I-CompBench [14]).
@@
The prompt set in GeckoNum is vastly more comprehensive and it systematically covers various dimensions of evaluation that affect numerical reasoning, such as different number ranges, types of nouns, different ways of representing numbers, approximations based on linguistic quantifiers, and reasoning about partial quantities that are missing in other datasets. We show that to evaluate a specific capability thoroughly requires an extensive, comprehensive dataset. Most similar to our work is [19], which investigates number generation for text-to-image models. We go beyond this by evaluating a comprehensive set of model families, considering other properties of numerical reasoning such as estimation and conceptual quantitative reasoning, and a more comprehensive breakdown of performance based on how commonplace an entity is.
@@ 3.1 Task 1: Exact Number Generation
To probe how well models capture The abstraction principle, we vary the context of numerical terms appear in. Then, we investigate whether and how image generation accuracy changes depending on the prompt structure (e.g., attributes, and compositionality) enumerated below to give seven different prompt types listed together with examples in rows “Exact” in Table 1.
Prompt structure: The simplest setting consists of phrases with an object and a number (numeric-simple). We explore different prompt structure of sentences (numeric-sentence), and also consider prompts which contain two or three number-noun combinations (2-additive and 3-additive). We then include color adjectives for both simple prompts with one entity (attribute-color), and two entities with associated colors (2-additive-color). Finally, we combine number terms with spatial relationships such as “There are three cats above one manatee.” (attribute-spatial).
Exact number generation: For the simplest setting (numeric-simple), we additionally examine the role of three other factors: (i) Number magnitude for which we generate text prompts with numbers ranging from 1 to 10. (ii) Number representation where we consider prompts that represent digits with both Arabic numerals (e.g., 1, 2, 3) or words (e.g., “one”, “two”, “three”). (iii) Noun frequency in English for which we select nouns (i.e., entities) to cover both frequent and rare words.111We use the wordfreq Python library [24] to determine word frequency. Our vocabulary in this task consists of approx. words from four common categories: food, animals, nature, and objects. See Appendix A for more details.
@@ 5 Evaluating Text-to-Image Models
Text-to-image models. We study seven different text-to-image models across three different model families: DALL·E 3 [2], Imagen-based models [21, 27], and Muse [4]. DALL·E 3 and Imagen models are pixel-based diffusion models, while the Muse family uses a different generative approach based on predicting masked image patches. Models within each family may differ in their size, architecture and training data. For models that have been trained on internal data sources we use letters of the alphabet to refer to earlier versions of models (i.e. Imagen-A is an earlier version of Imagen-B, and Imagen-D is the most recent version [27]). For each model we generate five images using five different seeds.
@@
| Random |
Overall performance. Table 2 shows average performance of each model on all tasks, with the baselines providing a reference on what would the accuracy be assuming an annotator randomly responded to the question. Overall, looking at how far model accuracies are from the baseline, we find that Task 1 is easiest for all models, while Task 3 is the hardest as performance of all models on it is close to or worse than that baseline. On Tasks 1 and 2, DALL·E 3 outperforms other models by a large margin, followed by Muse-B on Task 1 and Imagen-D on Task 2. However, even for Task 1, performance of the best model is still below 50%. We also find that more recent models do not necessarily perform better than earlier models (e.g. Imagen-D Imagen-C on Tasks 1 and 3).
Per prompt template breakdown of accuracy for each task in Figure 4 reveals further variability in difficulty, even within a single task.222For Task 1 we only use prompts with the same number range (i.e. 1–4) and vocabulary (words “apple”, “cat”, “koala”, “mushroom” and “bottle”) to dissociate the effect of the prompt template from other factors. For example, within Task 1, when we added another number term in the text prompt (c.f. numeric-simple and 2-additive rows in Figure 4) the accuracy on exact number generation for all models dropped substantially. One of the most difficult prompt categories in this task was attribute-spatial. In Task 3, the fractional-simple prompts appear easier than the fractional-complex prompts.
Finally, we find that while DALL·E 3 is the overall best performing model, there are certain cases where other models perform better. For example, Muse-B appears stronger at additive prompts (e.g. 2-additive, 2-additive-color and 3-additive), while Imagen-C appears stronger at numeric-simple. These models might indeed generate images that are better aligned with those prompt types, or they might be more accurate at the specific words present in the prompts under question. In the following sections, we further dissociate the effect of such variables while focusing on the strongest models. Results for all models, including the results of statistical significance tests used for model comparisons that follow are included in Appendix B.333Unless otherwise specified, we use the significance level when rejecting the null hypothesis based on the p-value of the Chi-squared test.
@@ 5.1 Task 1: Exact Number Generation
Number magnitude. We find that number magnitude strongly affects model performance in number generation, consistent with existing work in the multimodal and generative models [19, 18]. Figure 4 shows normalized confusion matrices for DALL·E 3 (A) and Imagen-A (B). The accuracy of the strongest model (DALL·E 3) drops substantially with each successive number (18 p.p. decrease for 1 2, 9 p.p. for 2 3, and 23 p.p. for 3 4). All models have a tendency to overestimate numbers (i.e. they tend to depict a higher number of entities than what is specified in the prompt), as shown by the non-zero entries over the “10+” label on the x-axis in the figure. With the exception of DALL·E 3, we also see underestimation, including instances where models fail entirely to generate the entity in the prompt (non-zero entries over the “0” label).
@@
Number representation (digits vs. words). Based on numeric-simple prompts and the small number range where models have higher accuracy (1–4), we find that six out of seven models, with DALL·E 3 being the exception, are significantly more accurate when numbers in text prompts are represented with words as opposed to digits. Figure 4C) shows accuracy for a subset of best performing models in each family for Task 1.
@@
Noun frequency (frequent vs. rare words). Within numeric-simple prompts type, we find significant differences in accuracy between prompts that include frequent nouns and prompts that include rare nouns for five out of seven models, with DALL·E 3 and Imagen-D being the exceptions. A subset of models is shown in Figure 4Ḋ). While this finding may indicate that these two models are better at abstracting the notion of a number as opposed to memorizing frequently occurring number-noun combinations, we note that this finding is based on a smaller subset of words that may or may not generalize for a greater sample size.
@@
Prompt structure. Prompts types with more than one number-noun combination, such as 2-additive and 3-additive, were on average more difficult compared to prompts with a single noun and a number (i.e. numeric-simple) (c.f. Table 4). Specifically, when numbers 2+ co-occurred with other numbers in the prompt, the accuracy of all models on these numbers dropped substantially compared to when those numbers appeared standalone, and the drop was larger for 3-additive than for 2-additive (see Section B.1.3 in the Appendix). With the exception of DALL·E 3, performance of all models was also significantly lower when numbers appeared in the same context with spatial relationships as in attribute-spatial prompts, as those prompts were significantly harder when compared to 2-additive prompts that did not include spatial relationships. Adding color terms to numbers significantly reduced accuracy only in some cases for some Imagen models and Muse-B. These results indicate that prompt complexity, such as including additional prepositions or numbers in the text, can dramatically impair the ability of models to correctly generate images containing even a small numbers of entities. In Section B.2 in the Appendix we discuss different types of model failures in Task 1, and in Section C.3.1 we discuss the occasional cases of disagreements in human annotations with several qualitative examples.
@@ 5.2 Task 2: Approximate Number Generation and Zero
When testing for approximate number generation, similarly to Task 1, we observe that prompts with fewer entities (i.e. approx-1-entity) are on average easier than prompts with two entities (i.e. approx-2-entity), as seen in Figure 4. Figure 5 shows a further breakdown of accuracy per specific prompt template. For approx-1-entity (first 3 rows containing only “X”), all seven models were the most accurate in generating “many” objects, and all seven models were least accurate in generating images with zero objects (i.e. “There is/are no X.”) For approx-2-entity (last 3 rows, with both “X” and “Y”), eight out of 10 models have highest accuracy when generating images from prompts of the “more X than Y”-type and lowest accuracy when generating “as many X as Y”. Our results highlight that understanding of linguistic quantifiers, even as simple as “no” is still challenging for generative models. Section C.3.2 in the Appendix shows qualitative examples including images and annotations in Task 2.
@@ 5.3 Task 3: Conceptual Quantitative Reasoning
Task 3 is the hardest, as most models perform close to or below the baseline (c.f. Table 2). Six out of seven models had the highest accuracy on fractional-simple prompts, followed by part-whole and fractional-complex, consistent with intuitions regarding the prompt difficulty. For DALL·E 3 differences in accuracies between the three prompt types were not significant. The last column in Figure 1 shows images for a prompt in the fractional-simple category, where many models failed to correctly depict “one apple and half of another apple”. Further examples, including images and annotations, are shown in Table A15.
However, evaluating models on Task 3 prompts was more challenging due to the lack of clear ground truth as in Task 1 (number in the prompt) and Task 2 (linguistic quantifier)—for some prompts, the DSG method generates only two relevant questions out of which one asks about the presence of an object in the image (e.g. Are there apples?) and thus answering “yes” yields the baseline of 50%, without necessarily capturing the nuance between whole apples, sliced apples or thirds/quarters of an apple. Further methodological challenges related to evaluation of models on this task are discussed in Appendix C.4.
@@ 6 Measuring What Counts: Challenges in Evaluation of Numerical Reasoning
Evaluating auto-eval metrics. Developing metrics that can reliably replace human evaluation is an active area of research. In particular, a recent line of work proposes auto-eval metrics for measuring text-to-image alignment by using pretrained language and/or VLMs [12, 13, 29, 7, 31]. Here, we evaluate such auto-eval metrics on their ability to capture exact number generation by examining how well they distinguish between correctly and incorrectly generated images for prompts in Task 1.
We consider five different auto-eval metrics: CLIPscore [12], a metric based on computing similarity between the text and an image, question-answering (QA) based metrics such as TIFA [13], Gecko [29], DSG [7], as well as VNLI [31], a metric fine-tuned to predict alignment between images and text. We divide images generated by Imagen-C, Muse-B, and DALL·E 3 into two groups each based on whether the image was generated correctly, and test whether the distributions of scores in the two groups is identical. Results in Table 3 show that only VNLI was able to reliably distinguish between correct and incorrect image generations, while all QA-based methods were able to do so only some models.
We also investigated whether auto-metrics are able to rank models based on accuracy. Following the approach in [29], we use a Wilcoxon signed-rank test to test whether there is a significant difference in average scores between models for the same metric, for all three combinations of model pairs. If such a difference exists, models are ranked based on the average score. We find that only Gecko, TIFA, and VNLI reliably rank models based on model accuracy (i.e. DALL·E 3 > Muse-B > Imagen-C). Overall, VNLI seems to be the most reliable auto-metric in evaluation of numerical reasoning.
@@ 7 Discussion and Conclusion
In this work, we propose GeckoNum, a benchmark that specifically focuses on measuring numerical reasoning capacity through three tasks: (i) exact, (ii) approximate number generation, and (iii) conceptual reasoning about quantities. We define various prompt types within these tasks to better control for the effect of factors such as the context of the number in the sentence and the sentence structure. We use GeckoNum to evaluate seven different text-to-image models from three different model families (DALL·E 3, Imagen, and Muse) by collecting human annotations for the generated images. We find that DALL·E 3 has the highest overall accuracy on exact and approximate number generation, while being the least impacted by some prompt manipulations we investigate (e.g. numbers represented as words vs. digits, frequent vs rare words); however, its performance on these tasks is still close to or under 50%. Depiction of approximate quantities and the concept of zero was also a weak point in all models. For example, even DALL·E 3 would consistently fail to correctly generate “a watermelon with no seeds” or “a cake with no candles”. We also find that the task requiring reasoning about parts and fractions was challenging for all models as their performance was close to the baseline.
@@
We used GeckoNum to highlight that many auto-eval metrics do not reliably differentiate between correct and incorrect images on numerical prompts—among metrics we tested, our preliminary experiments show that VNLI [31] is the only auto-eval metric capable of such distinction. Our results may suggest that current auto-eval metrics may require a specific adaptation for evaluation of numerical reasoning, either through fine-tuning, as done for VNLI, or by incorporating number-focused methods that require object detection [8]. We also demonstrated that our benchmark could be used to study and potentially improve the performance of pretrained vision–language models on counting.
@@
Overall, our results indicate that the current text-to-image models do not form abstract representations of numbers as their ability to reason about numbers is rudimentary: it is limited to depiction of exact, small quantities in images, and models frequently fail to generate approximate quantities and zero. We highlight the effectiveness of numerical reasoning as an open challenge in evaluation, since even in the large-scale training regime, due to the large combinatorial space of numbers, it is challenging to mitigate models’ weaknesses with more training data. Future modeling innovations—and not only better training data—might be needed to improve model performance on numerical reasoning. Our results show that most models are sensitive to the number representation in a given prompt (words vs digits). This showcases another interesting challenge involved in evaluating for numbers: they can occur in various formats (e.g., dates, phone numbers, or expressions such as “4k”) and representations. Also, effective ways of tokenizing numbers, again due to the vast space of possible numbers, is an open problem in language modeling [23] and an interesting direction for future research.

### v2

@@ 1 Introduction
Using GeckoNum, we evaluate twelve models chosen from five different model families (DALL·E 3, Midjourney, Imagen, Muse and Stable Diffusion): we generate images for these models and collect human annotations to measure whether the images correctly match the prompts with respect to numerical reasoning. Our benchmark consists of text prompts, generated images, and a total of human annotations that we release.111 https://github.com/google-deepmind/geckonum_benchmark_t2i
@@ 2 Related Work
Text-to-image benchmarking. Benchmarks used for evaluation of text-to-image models typically consist of a set of text prompts that target a specific capability.
Some benchmarks, such as DALL·E 3 Eval [2] are more general and aim to capture real world use cases, while others contain a variety of challenges, such as prompts in DrawBench [25], PartiPrompts [36], and HEIM [17]. Others are developed to interrogate models on more specific capabilities such as alignment (e.g. TIFA [15], DSG [9] and Gecko [33]) or compositionality (e.g. T2I-CompBench [16]).
@@
The prompt set in GeckoNum is vastly more comprehensive and it systematically covers various dimensions of evaluation that affect numerical reasoning, such as different number ranges, types of nouns, different ways of representing numbers, approximations based on linguistic quantifiers, and reasoning about partial quantities that are missing in other datasets. We show that to evaluate a specific capability thoroughly requires an extensive, comprehensive dataset. Most similar to our work is [22], which investigates number generation for text-to-image models. While [31] show that number generation improves with larger model sizes, their set of prompts is small (N=59) and focuses only on the simplest prompt structure. We go beyond this by evaluating a comprehensive set of model families, considering other properties of numerical reasoning such as estimation and conceptual quantitative reasoning, and a more comprehensive breakdown of performance based on how commonplace an entity is.
@@ 3.1 Task 1: Exact Number Generation
To probe how well models capture The abstraction principle, we vary the context of numerical terms appear in. Then, we investigate whether and how image generation accuracy changes depending on the prompt structure (e.g., attributes, and compositionality) enumerated below to give seven different prompt types listed together with examples in rows “Exact” in Table 1. Prompt structure: The simplest setting consists of phrases with an object and a number (numeric-simple). We explore different prompt structure of sentences (numeric-sentence), and also consider prompts which contain two or three number-noun combinations (2-additive and 3-additive). We then include color adjectives for both simple prompts with one entity (attribute-color), and two entities with associated colors (2-additive-color). Finally, we combine number terms with spatial relationships such as “There are three cats above one manatee.” (attribute-spatial).
Exact number generation: For the simplest setting (numeric-simple), we additionally examine the role of three other factors: (i) Number magnitude for which we generate text prompts with numbers ranging from 1 to 10. (ii) Number representation where we consider prompts that represent digits with both Arabic numerals (e.g., 1, 2, 3) or words (e.g., “one”, “two”, “three”). (iii) Noun frequency in English for which we select nouns (i.e., entities) to cover both frequent and rare words.222We use the wordfreq Python library [28] to determine word frequency. Our vocabulary in this task consists of approx. words from four common categories: food, animals, nature, and objects. See Appendix A for more details.
@@ 5 Evaluating Text-to-Image Models
Text-to-image models. We study 12 different text-to-image models across five different model families: DALL·E 3 [2], Midjourney, Imagen-based models [25, 31], Muse [4] and Stable Diffusion (SD) models [21, 23, 24]. The models we evaluate cover a wide span of model architectures, including pixel-based (e.g. Imagen) and latent-based (e.g. Muse, SD) models. Specifically, the Muse family uses a different generative approach based on predicting masked image patches.
Models within each family may differ in their size, architecture and training data. For models that have been trained on internal data sources we use letters of the alphabet to refer to earlier versions of models (i.e. Imagen-A is an earlier version of Imagen-B, and Imagen-D is a recent version referred to as Vermeer [31]). Generally, the core components of earlier Imagen models have fewer parameters compared to later models (e.g. 600M for Imagen-A, 2B for Imagen-B and Imagen-C, and 8B for Imagen-D). Imagen A, B and C models have been trained on internally curated versions of WebLI dataset [7], while Imagen-D has been trained on CC12M [5]. In all our experiments, for each model we generate five images using five different seeds.
@@
| SD1.5 | |||
| SD2.1 | |||
| SDXL | |||
| SD3 | |||
| Random Chance |
@@ 5.1 Task 1: Exact Number Generation
Number magnitude. We find that number magnitude strongly affects model performance in number generation, consistent with existing work in the multimodal and generative models [22, 20]. Figure 4 shows normalized confusion matrices for DALL·E 3 and Midjourney v6, the two strongest models on this task. Even for DALL·E 3, the accuracy drops substantially with each successive number (18 p.p. decrease for 1 2, 9 p.p. for 2 3, and 23 p.p. for 3 4). All models have a tendency to overestimate numbers (i.e. they tend to depict a higher number of entities than what is specified in the prompt), as shown by the non-zero entries over the “10+” label on the x-axis in the figure. With the exception of DALL·E 3, we also see underestimation, including instances where models fail entirely to generate the entity in the prompt (non-zero entries over the “0” label).
@@
Number representation (digits vs. words). Based on numeric-simple prompts and the small number range where models have higher accuracy (1–4), we find that 10 out of 12 models are significantly more accurate when numbers in text prompts are represented with words as opposed to digits, with the exception of DALL·E 3 and SD1.5, where there was no significant difference. Figure 4C) shows accuracy for a subset of best performing models in each family for Task 1.
@@
Noun frequency (frequent vs. rare words). Nine out of 12 models were more accurate when the nouns associated with numbers in text prompts were frequent words, as opposed to rare words, for numeric-simple prompts. The only exceptions were DALL·E 3, Imagen-D and SDXL, where there was no significant difference in accuracy. While this finding may indicate that these three models are better at abstracting the notion of a number as opposed to memorizing frequently occurring number-noun combinations, we note that this finding is based on a small set of words that may or may not generalize for a greater sample size. The comparison of accuracies for a subset of models is shown in Figure 4Ḋ).
@@
Prompt structure. Compositional prompts types with more than one number-noun combination, such as 2-additive and 3-additive, were on average more difficult compared to prompts with a single noun and a number (i.e. numeric-simple) (c.f. Figure 4). With the exception of DALL·E 3, performance of all other eleven models was also significantly lower when numbers appeared in the same context with spatial relationships as in attribute-spatial prompts, as those prompts were significantly harder when compared to 2-additive prompts that did not include spatial relationships. Adding color terms to numbers significantly reduced accuracy in Imagen-C, SD1.5 and SD3 models. These results indicate that prompt complexity, such as including additional prepositions or numbers in the text, can dramatically impair the ability of models to correctly generate images containing even a small numbers of entities. In Section B.2 in the Appendix we discuss different types of model failures in Task 1, and in Section C.3.1 we discuss the occasional cases of disagreements in human annotations with several qualitative examples.
@@ 5.2 Task 2: Approximate Number Generation and Zero
When testing for approximate number generation, similarly to Task 1, we observe that prompts with fewer entities (i.e. approx-1-entity) are on average easier than prompts with two entities (i.e. approx-2-entity), as seen in Figure 4. Figure 5 shows a further breakdown of accuracy per specific prompt template. For approx-1-entity (first 3 rows containing only “X”), Eleven out of 12 models were the most accurate in generating “many” objects, and all models were least accurate in generating images with zero objects (i.e. “There is/are no X”). For approx-2-entity (last 3 rows, with both “X” and “Y”), 8 out of 12 models have highest accuracy when generating images from prompts of the “more X than Y”-type and lowest accuracy when generating “as many X as Y”. Our results highlight that understanding of linguistic quantifiers, even as simple as the word “no”, can still be challenging for generative models. Section C.3.2 in the Appendix shows qualitative examples including images and annotations in Task 2.
@@ 5.3 Task 3: Conceptual Quantitative Reasoning
Task 3 is the hardest, as most models perform close to or below the random chance baseline (c.f. Table 2). All twelve models had the highest accuracy on fractional-simple prompts, followed by part-whole and fractional-complex for 10 out of 12 models, consistent with intuitions regarding the prompt difficulty. The last column in Figure 1 shows images for a prompt in the fractional-simple category, where many models failed to correctly depict “one apple and half of another apple”. Further examples, including images and annotations, are shown in Table A15.
However, evaluating models on Task 3 prompts was more challenging due to the question-based method used—for some prompts, the DSG method generates only two relevant questions out of which one asks about the presence of an object in the image (e.g. Are there apples?) and thus answering “yes” yields the baseline of 50%, without necessarily capturing the nuance between whole apples, sliced apples or thirds/quarters of an apple. Further methodological challenges related to evaluation of models on this task are discussed in Appendix C.4.
@@ 6 Measuring What Counts: Challenges in Evaluation of Numerical Reasoning
Evaluating auto-eval metrics. Developing metrics that can reliably replace human evaluation is an active area of research. In particular, a recent line of work proposes auto-eval metrics for measuring text-to-image alignment by using pretrained language and/or VLMs [14, 15, 33, 9, 35]. Here, we evaluate such auto-eval metrics on their ability to capture exact number generation by examining how well they distinguish between correctly and incorrectly generated images for numeric-simple prompts containing small numbers (i.e. 1–4) in Task 1.
We consider five different auto-eval metrics: CLIPscore [14], a metric based on computing similarity between the text and an image, question-answering (QA) based metrics such as TIFA [15], Gecko [33], DSG [9], as well as VNLI [35], a metric fine-tuned to predict alignment between images and text. We divide images generated by DALL·E 3, Midjourney v6, Imagen-C, Muse-B, and SD3 into two groups each based on whether the image was generated correctly or not. Then, we statistically test whether the distributions of scores in the two groups is identical. If there is a statistically significant difference between the two distributions, and if the correct generations have a higher score, we mark that case with a in Table 3. Gecko, DSG and VNLI are able to reliably distinguish between correct and incorrect image generations for all models. Upon manual inspection of generated questions for each QA method, we observed that Gecko and DSG generate similar questions, while TIFA occasionally generates questions about concepts that do not exist in the prompt. Following the approach in [33], we use a Wilcoxon signed-rank test to test whether there is a significant difference in average scores between pairs of models for the same metric. Only VNLI and Gecko were able to correctly order models in pairwise comparisons, in 8/10 and 7/10 cases, respectively. We also note that this analysis is based on simple prompts which trigger simple questions and thus most of the heavy lifting in performance of an auto-metric is delegated to the underlying VQA model.
@@ 7 Discussion and Conclusion
In this work, we propose GeckoNum, a benchmark that specifically focuses on measuring numerical reasoning capacity through three tasks: (i) exact and (ii) approximate number generation, and (iii) conceptual reasoning about quantities. We define various prompt types within these tasks to better control for the effect of factors such as the context of the number in the sentence and the sentence structure. We use GeckoNum to evaluate twelve different text-to-image models from five different model families (DALL·E 3, Midjourney v6, Imagen, Muse and Stable Diffusion) by collecting human annotations for the generated images. We find that DALL·E 3 has the highest overall accuracy on exact and approximate number generation, while being the least impacted by some prompt manipulations we investigate (e.g. numbers represented as words vs. digits, frequent vs rare words); however, its performance on these tasks is still close to or under 50%. Depiction of approximate quantities and the concept of zero was also a weak point in all models. For example, even DALL·E 3 would consistently fail to correctly generate “a watermelon with no seeds” or “a cake with no candles”. We also find that the task requiring reasoning about parts and fractions was challenging for all models as their performance was close to the baseline.
@@
We used GeckoNum to show that only some auto-eval metrics can reliably differentiate between correct and incorrect images on simple numerical prompts—among metrics we tested, our preliminary experiments show that VNLI [35] and Gecko [33] are the only two metrics capable of such differentiation, including the ability to rank models in pairwise model comparisons. We also demonstrated that our benchmark could be used to study and potentially improve the performance of pretrained vision–language models on counting.
@@
Our approach relies on human annotators to provide counts of objects in an image, as is often the gold standard in evaluation of text-to-image models. However, such annotation process is laborious, costly and does not scale. We expect that the rapid improvement of pipelines and frameworks for evaluation of text-image alignment will reduce the need for manual annotation [10, 35, 13]. In addition, we identified three important numerical capabilities for which we have manually designed human annotation templates, but numerical cognition in humans spans a wide set of gradually developing capabilities that might require additional evaluation templates. As discussed in Section 5.3, the design of protocols for evaluation of more complex aspects of numerical reasoning remains an important open challenge with rapid advancements in models. Overall, our results indicate that the current text-to-image models do not form abstract representations of numbers as their ability to reason about numbers is rudimentary: it is limited to depiction of exact, small quantities in images, and models frequently fail to generate approximate quantities and zero. We highlight the effectiveness of numerical reasoning as an open challenge in evaluation, since even in the large-scale training regime, due to the large combinatorial space of numbers, it is challenging to mitigate models’ weaknesses with more training data. Future modeling innovations—and not only better training data—might be needed to improve model performance on numerical reasoning. Our results show that most models are sensitive to the number representation in a given prompt (words vs digits). This showcases another interesting challenge involved in evaluating for numbers: they can occur in various formats (e.g., dates, phone numbers, or expressions such as “4k”) and representations. Also, effective ways of tokenizing numbers, again due to the vast space of possible numbers, is an open problem in language modeling [27] and an interesting direction for future research.